OpenStudy (anonymous):

PV = nRT Ideal gas law There is a gas in a nonrigid container. If "n" is halved and T remains always constant, how will P and V change in response? Is it possible to make a conclusion given only this information?

OpenStudy (aaronq):

you can go about it in a mathematical way: since you have T held constant you can omit it (also R is just a constant). so you end up with PV=n suppose you have a system where they all equal 1, now suppose you reduce the number of moles, n, to 1/2 (one half) solving for P, P(1)=1/2 -> P=1/2, so the pressure decreased proportional to the number of moles. similarly if you solve for V, (1)V=1/2, V=1/2, the volume (given that the container is flexible, like a ballon) would decrease. Now note that if you say lower the number of moles and then either the volume or pressure decreases, the other could remain constant. eg. say if you have a balloon with x amount of gas, if you let some gas out, the volume will decrease, but the pressure COULD remain constant or increase or decrease.